xdy + (y - x2 y2 ) dx = 0


1
Expert's answer
2021-07-20T15:53:42-0400

Solution

"xdy + (y - x^2y^2 ) dx = 0"


Divide through by "dx" as follows

"x{dy\over dx} + (y - x^2y^2 ){ dx\over dx} = {0\over dx}"

We get

"x{dy\over dx} + (y - x^2y^2 ) =0"

"x{dy\over dx} + y - x^2y^2 =0"

This is a first-order nonlinear differential equation

Let "u=xy"

Use product rule of differentiation to differentiate "u=xy" with respect to "x"

"u=xy"

"{du\over dx}=({d\over dx}(x)).y+x.({d\over dx}(y))"

"\implies {du\over dx}=y+x{dy\over dx}"

From "{du\over dx}=y+x{dy\over dx}" ,We can get expression for "{dy\over dx}"

"{du\over dx}=y+x{dy\over dx}"

"x{dy\over dx}={du\over dx}-y"

Now, divide through by "x" to get

"{dy\over dx}={{du\over dx}-y \over x}={{du\over dx}\over x}-{y\over x}"


From "u=xy" , "y={u\over x}"


"\therefore {dy\over dx}={{du\over dx}\over x}-{{u\over x}\over x} ={{du\over dx}\over x}-{u\over x^2}"


Now, let's substitute "{dy\over dx}" and "y" to the differential equation


"x{dy\over dx} + y - x^2y^2 =0"


"x({{du\over dx}\over x}-{u\over x^2})+ {u\over x} - x^2({u\over x})^2 =0"


"{{du\over dx}}-{u\over x}+ {u\over x} - x^2.{u^2\over x^2} =0"


"{{du\over dx}}- u^2=0"


"\implies {du\over dx}=u^2 \implies {du\over u^2}=dx"

Take integral on both sides of "{du\over u^2}=dx"


we have


"\int{du\over u^2}=\int dx \implies -{1\over u}=x+C \implies u=-{1\over x+C}"


Recall "y={u\over x} \implies y={-{1\over x+C}\over x}=-{1\over x+C}\times{1\over x}=-{1\over x(x+C)}"


"\therefore" the general solution is thus "y=-{1\over x(x+C)}"